OpenStudy (anonymous):

If on an average 1 ship in every 10 sinks,find the chance that out of 5 ships atleast 4 will arrive safely.

OpenStudy (anonymous):

Attempt: Net P=P(4)+P(5) Using Bernoulli trials> \[\Large C(5,4) (1-\frac{1}{10})^4 (\frac{1}{10})^1+C(5,5) (1-\frac{1}{10})^5\]

OpenStudy (anonymous):

\[\LARGE 5 \times (\frac{9}{10})^4 \times \frac{1}{2}+ (\frac{9}{10})^5\]

OpenStudy (anonymous):

\[\LARGE \frac{9^4}{10^5}(5+9)= \frac{9^4}{10^5} \times 14\]

OpenStudy (anonymous):

=0.98

OpenStudy (anonymous):

answer is 0.9185

OpenStudy (anonymous):

oops a little mistake

OpenStudy (anonymous):

\[\LARGE (\frac{9}{10})^4 \times \frac{1}{2}+ (\frac{9}{10})^5\]

OpenStudy (anonymous):

\[\LARGE \frac{9^4}{10^4}(1+0.9)=\frac{9^4}{10^4}(1.9)\]

OpenStudy (anonymous):

=1.24 o.O

OpenStudy (anonymous):

duh! silly mistakes got it thanks